OpenStudy (anonymous):

Tutorial On Finding General Solution for Trigonometric Equations:

OpenStudy (anonymous):

General Solution: Normally, if we are asked to find the solution for \(\theta\) in the interval \((0, 2 \pi)\) for the equation : \(\large \color{red}{\cos(\theta) = \frac{1}{2}}\) then we say \(\large \theta = 60^{\circ} = \frac{\pi}{3}\), but in that interval this is not the only value of \(\theta\), \(\theta\) can have more values like : \(\large \theta = 300^{\circ} = \frac{5 \pi}{3}\).. As all the trigonometric ratios are cyclic or circular in nature, so we will get infinite values for \(\theta\) that will satisfy the equation \(\large cos(\theta) = \frac{1}{2}\).. This we can achieve by Finding the General Solution for the given Trigonometric Equation.. -------------------------------------------------------- First Form: 1. For \(\color{blue}{sin(\theta) = 0} \) , the general solution for \(\theta\) is : \(\color{blue}{ \theta = n \pi }\) , where \(n \in Z (Integers)\).. Here, \(n\) can take integer values like -2, -1, 0, 1 , 2 etc.. 2. For \(\color{red}{cos(\theta) = 0} \) , the general solution for \(\theta\) is : \(\color{red}{\theta = (2n + 1) \frac{\pi}{2}}\), where \(n \in Z (Integers)\).. 3. \(\color{green}{tan(\theta) = 0 }\) , the general solution for \(\theta\) is : \(\color{green}{\theta = n \pi}\), where \(n \in Z (Integers)\).. Note: For \(sin(\theta) = 0\) and \(tan(\theta) = 0\) , the general solution is same.. Here is an example: 1. For \(\color{red}{cos(3 \theta) = 0}\) find the general solution.. Solution: \(cos(3\theta) = 0\) \(\implies 3 \theta = (2n + 1) \frac{\pi}{2}\) where \(n \in Z\) \(\implies \theta = (2n + 1) \frac{\pi}{6}\) where \(n \in Z\) ----------------------------------------------------------- Second Form: 1. For \(\color{green}{sin(\theta) = sin(\alpha)}\), the general solution is : \(\color{green}{\theta = n \pi + (-1)^{n} \alpha, n \in Z }\) 2. For \(\color{red}{cos(\theta) = cos(\alpha)}\), the general solution is : \(\color{red}{\theta = 2n \pi \pm \alpha, n \in Z }\) 3. For \(\color{blue}{tan(\theta) = tan(\alpha)}\), the general solution is : \(\color{blue}{\theta = n \pi + \alpha, n \in Z }\).. Here is an example: For \(\color{blue}{cosec(\theta) = 2}\) find the general solution.. Solution: \(cosec(\theta) = 2\) \(\implies \large \frac{1}{sin(\theta)} = 2\) or \(sin(\theta) = \frac{1}{2} \implies sin(\theta) = sin(\frac{\pi}{6})\) So: \(\theta = n \pi + (-1)^{n} (\frac{\pi}{6}) , n \in Z\) ------------------------------------------------------------- Third Form: 1. For \(\color{blue}{sin^2(\theta) = sin^2(\alpha)}\) and \(\color{blue}{cos^2(\theta) = cos^2(\alpha)}\) and \(\color{blue}{tan^2(\theta) = tan^2(\alpha)}\), the general solution for the three is same and given by: \(\color{blue}{\theta = n \pi \pm \alpha}\) where \(n \in Z\).. Here is an example for this: For \(\color{orange}{7cos^2(\theta) + 3 sin^2(\theta) = 4}\) find the general solution.. Solution: \[\color{red}{ 7cos^2(\theta) + 3 sin^2(\theta) = 4}\] \(\color{blue}{7(1 - sin^2(\theta)) + 3 sin^2(\theta) = 4 \implies 4sin^2(\theta) = 3 \implies sin^2(\theta) = \frac{3}{4}} \) \(\color{red}{\implies sin^2(\theta) = (\frac{\sqrt{3}}{2})^2 \implies sin^2(\theta) = sin^2 (\frac{\pi}{3})}\) \(\implies \large \color{green}{\theta = n \pi \pm \frac{\pi}{3}}\) where \(n \in Z\).. Like wise you can find the general Solution for all the Trigonometric Equations..

OpenStudy (anonymous):

i must say good tutorial. and great work. good for students have difficulties in finding the trigonometric values correctly. also good latex work:) you are good in the equation editing tool !

OpenStudy (anonymous):

Thanks @sami-21 ..

OpenStudy (anonymous):

Some people have curly brown hair,they paint black some=sine people have=perpendicular(from people)/hypatenus(from have) so sine=perpendicular/hypatanus curly=cosine brown hair=base(from brown)/hypataneus(from hair) so cosine=base/hypatenus they=tan paint black=perpendicular(from paint)/base(from black) so tan=perpendicular/base whole trignometry in just in line!!!!!!!!!!

OpenStudy (anonymous):

@sami-21 check it:))))

OpenStudy (anonymous):

i think u also know that @sami-21

OpenStudy (anonymous):

\[\huge \text{Nice} \]

OpenStudy (anonymous):

\[\huge \color{pink}{Like \quad you...}\]

OpenStudy (anonymous):

http://www.interactivetriangulation.com/T/001.html

OpenStudy (anonymous):

Let me try something . m quite new to this latex world :P \[\Huge \color{red} {Awesome!}\]

mathslover (mathslover):

\[\huge{\mathbb{G}\textbf{r}\mathbb{E}\textbf{a}\mathbb{T}\space \mathbb{W}\textbf{o}\mathbb{R}\textbf{k}}\mathbb{!}\]

OpenStudy (vishweshshrimali5):

good work .. @waterineyes .. thanks @mathslover for getting me noticed about such a good tutorial

OpenStudy (anonymous):

Thanks @vishweshshrimali5 and @mathslover

mathslover (mathslover):

-_--_-- _____ |-_-| _|| ||_ :)

OpenStudy (anonymous):

\[\huge \color{green}{^{ \cdot}\smile ^{\; \cdot}}\]

OpenStudy (anonymous):

Great Work. God bless you :)

OpenStudy (anonymous):

@annas thanks buddy..

mathslover (mathslover):

\[\large{\color{green}{^\cdot}\color{blue}{-}\color{white}{^\cdot}}\]

mathslover (mathslover):

notice the secnd eye also :)

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

I am unable to notice @mathslover

OpenStudy (anonymous):

That is closed or not ??

mathslover (mathslover):

select that :)

OpenStudy (anonymous):

kush its one eyed smiley :P

mathslover (mathslover):

:D one green and other one is white .. notice that :)

OpenStudy (anonymous):

Yeah I noticed that.. Ha ha haha..

mathslover (mathslover):

:D

OpenStudy (anonymous):

Just like you.. Thanks @angela210793

OpenStudy (anonymous):

And me also..

OpenStudy (lgbasallote):

torture for the epileptics :C lol

OpenStudy (anonymous):

yes, very nice indeed. i especially like "second form" and "third form" now allow me one small criticism the line \(\large \theta = 60^{\circ} = \frac{\pi}{3}\) it is certainly true that the angle whose measure is 60 degrees is the same as the angle whose radian measure is \(\frac{\pi}{3}\) but it is not true that the numbers are the same. as an analogy, the temperature of 77 F is the same as the temperature of 25 C but you would not say \(77=25\) since as functions of numbers, the trig functions correspond to functions of angles in radians, not degrees, (which is why, for example, the derivative of sine is not cosine, if you are working in degrees) it is best to stick to numbers rather than angles.

OpenStudy (anonymous):

I will surely keep it in mind @satellite73 Thanks for your feedback...

rvc (rvc):

awesome

OpenStudy (anonymous):

Thanks..!!

Nnesha (nnesha):

gO_Od work!! @eta or @tears